What us the work done by a system when it absorbs 47 J of heat and gain 12 J of energy
1
Expert's answer
2020-10-02T14:19:19-0400

According to first law of thermodynamics

du = dq + dw ; where du = change in internal energy, dq = heat change ( +ve when system gain heat and -ve when system loose heat) , dw = work done by the system(-ve) or work done on the system (+ve)

Here dq = 47 J

du = 12 J

dw = du - dq = 12 - 47 = -35 J

So work done by the system as the sign appear is -ve, so system will work 35 J.